OpenStudy (anonymous):

Confusing percale question...? Which of the following could be an example of a function with a domain [a, infinity) and a range (-infinity, b] where a > 0 and b > 0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you have to make one up?

OpenStudy (misty1212):

we can do that if you like, but if you have choices then it will be easier to pick one

OpenStudy (anonymous):

Here are my choices http://i296.photobucket.com/albums/mm191/cholland2015/precal_zpsoptusrgt.png

OpenStudy (anonymous):

make a guess, i bet you are right!

OpenStudy (misty1212):

lol one of those is the one i was going to suggest

OpenStudy (misty1212):

you want the domain to be \([a,\infty)\) i as going to suggest something involving the square root of \(x-a\)

OpenStudy (anonymous):

I'm confused of where the infinity part comes in?

OpenStudy (anonymous):

that just means the domain is all numbers larger than \(a\) don't fret about the \(\infty\)

OpenStudy (anonymous):

do what @misty1212 said, pick one with \(\sqrt{x-a}\) since you cannot take the square root of a negative number that would mean \(x-a\geq0\) so \(x\geq a\)

OpenStudy (anonymous):

Okay I see... I think. Thank you guys both for your help.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

btw do you know why the range of \(-\sqrt{x-a}+b\) is \((-\infty,b]\)?